# Existence, Sharp Boundary Asymptotics, and Stochastic Optimal Control for Semilinear Elliptic Equations with Gradient-Dependent Terms and Singular Weights

### 08 Mar 2026 v1

@@ Abstract.
We establish existence, uniqueness, and precise boundary asymptotic behavior for large solutions to the semilinear elliptic equation
$\Delta u+b(x)h(|\nabla u|)+a(x)u=f(x)$
in a bounded strictly convex domain $\Omega\subset\mathbb{R}^{N}$. Here, $h$ is a strictly convex function with power-like growth $h(s)\sim s^{q}$ for $q\in(1,2]$, and the weights $a(x),b(x)$ exhibit prescribed singular growth near $\partial\Omega$. Our main contributions are threefold: (i) we prove existence and uniqueness via Perron’s method and derive sharp liminf and limsup bounds for the blow-up rate $\gamma=(\beta-q+2)/(q-1)$, extending recent high-precision analytic techniques; (ii) we establish the strict convexity of solutions through the microscopic convexity principle; (iii) we provide a rigorous verification theorem identifying the solution as the value function of an infinite-horizon stochastic control problem with state constraints, in the spirit of the Lasry–Lions framework. Three distinct asymptotic regimes are identified based on the interplay between the gradient growth and the weight singularity. Numerical experiments using a monotone iterative scheme validate the theoretical boundary limits and the geometric properties of the solutions.
@@
Semilinear elliptic equations; boundary blow-up; exact large solutions; Hamilton–Jacobi–Bellman equations; stochastic optimal control; state constraints; viscosity solutions
@@ 1. Introduction
The study of boundary blow-up solutions (also known as large solutions) for elliptic equations has a celebrated history, dating back to the pioneering works of Bieberbach [6] and Rademacher [22] on the equation $\Delta u=f\left(u\right)$ with $f\left(u\right)=e^{u}$. These early investigations showed that the blow-up rate is determined by the interplay between the domain geometry and the nonlinearity structure. The modern theory was significantly advanced by Keller [14] and Osserman [21], who established the necessary and sufficient condition for existence: $\int_{1}^{\infty}[F(s)]^{-1/2}ds<\infty$, $F(s)=\int_{0}^{s}f\left(t\right)dt$. This Keller–Osserman condition has since been generalized to a wide class of operators and settings; see [20, 7] for comprehensive reviews.
1.1. State of the Art and Motivation
The study of boundary blow-up solutions (large solutions) for elliptic equations has a deep history, beginning with Bieberbach [6] and Rademacher [22]. The theory was formalized by Keller [14] and Osserman [21], who introduced the Keller–Osserman condition for existence. In recent decades, research has shifted towards more complex operators involving gradient terms and singular weights. Notable contributions by Marcus and Véron [20] and Covei [7, 8] have explored the interplay between domain geometry and nonlinearity.
@@ 1.1. State of the Art and Motivation
A modern frontier in the field is the connection to stochastic optimal control, as pioneered by Lasry and Lions [17]. In their framework, boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations serve as value functions for control problems with state constraints. This dual interpretation provides both physical intuition and rigorous regularity results. Recent advances by Zhang [23] in the context of Monge-Ampère equations have introduced high-precision analysis for singular weights, which we adapt here for the Laplacian setting.
@@ 1.2. Main Contributions and Organization
1.3. Assumptions
Throughout this paper, we adopt the following structural and geometric assumptions:
@@ 1.3. Assumptions
$\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded domain with $C^{2,\alpha}$ boundary for some $\alpha\in(0,1)$. Moreover, $\Omega$ is strictly convex: there exists $\kappa_{0}>0$ such that all principal curvatures of $\partial\Omega$ are at least $\kappa_{0}$.
@@
$f:\overline{\Omega}\to[0,\infty)$ is a Hölder continuous function.
@@
    $h$ is strictly increasing and strictly convex;
@@
    There exist constants $l_{1},l_{2}>0$ and $q\in(1,2]$ such that $l_{1}s^{q}\leq h(s)\leq l_{2}s^{q}$ for all $s\geq 0$.
@@
The weights $a,b\in C^{\infty}(\Omega)$ are strictly positive in $\Omega$ and satisfy:
@@
    $a_{1}d(x)^{\alpha}\leq a(x)\leq a_{2}d(x)^{\alpha},$ (1.2)
    near $\partial\Omega$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$.
@@
    There exists a constant $\beta\in\mathbb{R}$ and limits $b_{1},b_{2}>0$ such that
    $\liminf_{d(x)\rightarrow 0}\frac{b(x)}{d(x)^{\beta}}=b_{1}\leq\limsup_{d(x)\rightarrow 0}\frac{b(x)}{d(x)^{\beta}}=b_{2}.$ (1.3)
This structure on the weights perfectly mirrors the rigorous bounds employed by Zhang [23] in the context of Monge-Ampère equations with weights, enabling us to derive precise trace limits for our solutions.
1.4. The main results
Our main contributions are:
@@ 1.4. The main results
Under Assumptions 1.11.4, there exists a unique classical solution $u\in C^{2}(\Omega)$ to (1.1). Moreover, if $q<\beta+2$, this solution satisfies the blow-up condition and the asymptotic estimate
$C_{-}d(x)^{-\gamma}\leq u(x)\leq C_{+}d(x)^{-\gamma}\quad\text{near }\partial\Omega,$ (1.4)
where $\gamma=(\beta-q+2)/(q-1)$ and $C_{\pm}$ are explicit barrier constants.
@@
$\xi_{1}\leq\liminf_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}\leq\limsup_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}\leq\xi_{2},$ (1.5)
@@
If the limits $b_{0}=\lim_{d(x)\to 0}b(x)d(x)^{-\beta}$ and $l_{0}=\lim_{s\to\infty}h(s)s^{-q}$ exist, then:
$\lim_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}=\xi_{0}=\left(\frac{\gamma(\gamma+1)}{b_{0}l_{0}\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (1.7)
1.5. Original Contributions and Contextualization
This article bridges the gap between purely analytical boundary blow-up theory and the stochastic control framework of Lasry and Lions. Our primary original contributions are:
  1. (1)
    We prove existence and uniqueness for a class of semilinear equations on strictly convex (non-radial) domains with singular weights $a(x)$ and $b(x)$.
  2. (2)
    We derive exact sharp boundary asymptotics, isolating precise constants that depend on the singular behavior of the weights, mirroring the rigorous approach of Zhang [23].
  3. (3)
    We establish the strict convexity of solutions using the microscopic convexity principle, a result that holds special significance for the control-theoretic interpretation.
  4. (4)
    We provide a rigorous verification theorem and a convergent numerical algorithm that confirms the theoretical predictions.
1.6. Structure of the Paper
The remainder of the work is organized as follows. Section 3 provides the complete proof of Theorem 1.5, integrating barrier construction and Perron’s method. Section 4 details the proof of Theorem 1.6 regarding sharp boundary asymptotics. Section 5 explains the stochastic model of origin and the verification theorem. Section 6 presents the numerical analysis, including the monotone iterative scheme and visual results. Section 7 concludes the work, while Section References contains the alphabetical list of references.
1.7. Stochastic Control and the Lasry–Lions Framework
A fundamental connection between large solutions and stochastic optimal control was established in the seminal work of Lasry and Lions [17]. They demonstrated that boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations arise naturally as value functions for optimal control problems with state constraints. Specifically, if a controlled diffusion process is required to remain within a domain $\Omega$ for all time (a state constraint), then the associated value function satisfies an elliptic PDE with $u(x)\to+\infty$ as $x\to\partial\Omega$. The infinite boundary value acts as an “infinite penalty” that prevents the optimal trajectory from exiting the domain.
This stochastic control interpretation provides both physical motivation and powerful analytical tools. The theory has been further developed in the context of ergodic control [4], mean-field games [18], and viscosity solutions [10]; see also the article [8] and the monographs by Fleming and Soner [12] and Bensoussan and Lions [5].

### 22 Mar 2026 v2

@@ Abstract.
This article develops a rigorous general framework for a class of semilinear elliptic equations with gradient-dependent nonlinearities and singular weights posed in strictly convex domains. We study large solutions to
$\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u=f(x)\quad\text{in }\Omega\subset\mathbb{R}^{N},$
where $h$ is strictly convex with power-type growth $h(s)\sim s^{q}$ for $q\in(1,2]$, and the coefficients $a(x)$ and $b(x)$ exhibit prescribed singular behavior near $\partial\Omega$. The main contributions are fourfold. First, we establish existence and uniqueness of large solutions via Perron’s method and derive sharp boundary asymptotics, including optimal liminf–limsup estimates for the blow-up rate $\gamma=(\beta-q+2)/(q-1)$, thereby extending recent high-precision analytic techniques. Second, we prove the strict convexity of solutions by means of the microscopic convexity principle, revealing new geometric properties induced by the interplay between gradient growth and weight singularity. Third, we provide a verification theorem showing that the solution coincides with the value function of an infinite-horizon stochastic optimal control problem with non-violable state constraints, in the spirit of the Lasry–Lions approach. Fourth, we develop a comprehensive section on applications in Operational Research and Management Science, including real-world case studies in inventory management, portfolio risk control, and supply chain optimization, and we explore interdisciplinary connections with physics through boundary layer theory and diffusion processes. Numerical experiments for both the one-dimensional ($N=1$) and two-dimensional ($N=2$) cases, based on a monotone iterative scheme, confirm the theoretical boundary behavior and validate the predicted geometric structure. All Python implementation codes are provided in the Appendix.
@@
Semilinear elliptic equations; boundary blow-up; exact large solutions; Hamilton–Jacobi–Bellman equations; stochastic optimal control; state constraints; viscosity solutions; operational research; inventory management; portfolio optimization
@@ 1.1. Historical Background and State of the Art
The study of boundary blow-up solutions (also known as large solutions) for elliptic equations has a celebrated history, dating back to the pioneering works of Bieberbach [6] and Rademacher [24] on the equation $\Delta u=f(u)$ with $f(u)=e^{u}$. These early investigations showed that the blow-up rate is determined by the interplay between the domain geometry and the nonlinearity structure. The modern theory was significantly advanced by Keller [14] and Osserman [22], who established the necessary and sufficient condition for existence:
$\int_{1}^{\infty}[F(s)]^{-1/2}\,ds<\infty,\quad F(s)=\int_{0}^{s}f(t)\,dt.$ (1.1)
This Keller–Osserman condition has since been generalized to a wide class of operators and settings; see [20, 7] for comprehensive reviews.
In subsequent decades, the theory was extended in multiple directions. Marcus and Véron [20] developed a comprehensive framework for uniqueness and asymptotic behavior of solutions with boundary blow-up, establishing connections between the growth rate of the nonlinearity and the geometry of the domain. Díaz and Letelier [11] studied explosive solutions of quasilinear elliptic equations, introducing techniques that allow for more general differential operators. Bandle and Giarrusso [3] initiated the study of boundary blow-up for equations with nonlinear gradient terms, demonstrating that the interaction between the absorption term and the gradient term generates qualitatively new phenomena.
A pivotal development in the theory was the introduction of equations involving gradient-dependent nonlinearities. In such equations, the structure
$\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u=f(x)$ (1.2)
requires fundamentally different techniques from the classical Keller–Osserman framework, since the gradient term introduces a first-order singularity that competes with the second-order diffusion. The interplay between these terms determines the asymptotic behavior of the solution near the boundary, giving rise to distinct asymptotic regimes.
@@
A modern frontier in the field is the connection to stochastic optimal control, as pioneered by Lasry and Lions [17]. In their seminal 1989 paper, they demonstrated that boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations serve as value functions for control problems with state constraints. This dual interpretation provides both physical intuition and rigorous regularity results. Specifically, if a controlled diffusion process is required to remain within a domain $\Omega$ for all time (a state constraint), then the associated value function satisfies an elliptic PDE with $u(x)\to+\infty$ as $x\to\partial\Omega$. The infinite boundary value acts as an “infinite penalty” that prevents the optimal trajectory from exiting the domain.
This stochastic control interpretation has been further developed in the context of ergodic control [4], mean-field games [18], and viscosity solutions [10]; see also the article [8] and the monographs by Fleming and Soner [12] and Bensoussan and Lions [5]. Recent advances by Zhang [27] in the context of Monge-Ampère equations have introduced high-precision analysis for singular weights, which we adapt here for the Laplacian setting.
From the perspective of Operational Research, the connection between elliptic PDEs and stochastic control has far-reaching implications. The value function of a state-constrained stochastic control problem encodes the optimal cost-to-go for a manager who must keep a stochastic system within safe operating limits. This formulation arises naturally in inventory management [23, 28], portfolio optimization under risk constraints [21], and production planning in manufacturing systems [8, 26].
1.2. Limitations of Existing Models
Despite extensive research since the pioneering works of Bieberbach [6] and Rademacher [24], existing models frequently face limitations that hinder their general applicability. We identify four principal limitations:
Limitation 1: Geometric restrictions. Many classical studies are restricted to radial domains (balls), where the geometry reduces to a single ordinary differential equation. While these simplifications allow for elegant closed-form solutions, they fail to capture the complexity of real-world scenarios where the underlying state space is not symmetric. In economic applications, the feasible operating region is rarely a ball; it is typically a polyhedron, an ellipsoid, or a more general convex set defined by multiple constraints.
Limitation 2: Non-singular coefficients. The majority of existing results assume that the coefficients $a(x)$ and $b(x)$ are bounded and bounded away from zero. This assumption is incompatible with economic models where the cost of corrective action or the discount rate varies significantly with the state. For instance, in inventory management, the cost of emergency replenishment increases dramatically as stock approaches zero, corresponding to a weight $b(x)$ that blows up at the boundary.
Limitation 3: Absence of gradient terms. The Keller–Osserman condition (1.1) applies to equations of the form $\Delta u=f(u)$, where there is no gradient dependence. The introduction of a gradient term $h(|\nabla u|)$ fundamentally changes the problem structure, and the classical theory does not directly apply. A new balance equation must be derived that accounts for the competition between the Laplacian and the gradient term.
Limitation 4: Gap between theory and practice. Most existing results provide existence theorems under abstract conditions that are difficult to translate into operational control laws for managers or engineers. There is a significant gap between purely analytical investigations and their practical implementation. The theory must provide not only existence and uniqueness, but also explicit asymptotic formulas, convexity guarantees, and numerical algorithms that can be implemented in practice.
1.3. The New General Framework
The motivation for this work is the construction of a general theory for the class of singular problems described by (1.2). Our framework addresses all four limitations identified above:
  1. (i)
    We work on general bounded strictly convex domains $\Omega\subset\mathbb{R}^{N}$, not restricting to radial geometries.
  2. (ii)
    We allow the weights $a(x)$ and $b(x)$ to exhibit prescribed singular behavior near $\partial\Omega$, with explicit power-law rates.
  3. (iii)
    We handle general strictly convex Hamiltonians $h$ with power-type growth $h(s)\sim s^{q}$ for $q\in(1,2]$, fully accounting for the competition between diffusion and gradient terms.
  4. (iv)
    We provide explicit asymptotic constants, convexity results, and a stochastic control interpretation that directly yields optimal feedback control laws.
Our approach is built on the principle of independence from concrete data meaning. We argue that the fundamental dynamics of “preventive correction” are governed by universal geometric and analytical laws. This allows the researcher to apply our results to any system where a state $X_{t}$ must be kept in $\Omega$ by a drift $\xi_{t}$, provided the cost functional mirrors the singular structures we define. In this sense, the article provides general, high-precision tools in stochastic optimization that remain valid as long as the structural assumptions are met.
The key innovation in our model lies in the identification of three distinct asymptotic regimes determined by the relationship between the gradient exponent $q$ and the weight singularity $\beta$:
  • Gradient-Dominant Regime ($q<\beta+2$): The solution blows up as $u(x)\sim C\,d(x)^{-\gamma}$ with $\gamma=(\beta-q+2)/(q-1)>0$.
  • Critical Logarithmic Regime ($q=\beta+2$): The blow-up is logarithmic, $u(x)\sim C\,\ln(1/d(x))$.
  • High-Order Gradient Regime ($q>\beta+2$): The gradient term dominates at all orders.
This classification, combined with the exact analytical constants we derive, provides a complete picture of the boundary behavior. The universal correction constant $\xi_{0}$ that we identify represents the precise intensity of the boundary penalty required to enforce the state constraint, and it depends only on the structural parameters $(q,\beta,b_{0},l_{0})$ of the problem.
@@ 1.4. Main Contributions and Organization
Our main contributions are:
  1. (1)
    Existence and uniqueness (Theorem 2.9): We prove existence and uniqueness for a class of semilinear equations on strictly convex (non-radial) domains with singular weights $a(x)$ and $b(x)$, via Perron’s method and explicit barrier construction.
  2. (2)
    Sharp boundary asymptotics (Theorem 2.10): We identify three distinct regimes (gradient-dominant, high-order, and logarithmic) and derive the exact analytical constants that govern the blow-up rate, mirroring the rigorous approach of Zhang [27].
  3. (3)
    Convexity inheritance (Theorem 2.11): Under the additional assumptions that $a$ and $b$ are convex and $f$ is concave, we prove that the strict convexity of the boundary is inherited by the solution’s Hessian, ensuring the stability and uniqueness of optimal control strategies.
  4. (4)
    Verification theorem (Theorem 6.9): We prove that the analytical solution is the value function of a state-constrained stochastic control problem and derive the optimal feedback control law explicitly.
  5. (5)
    Economic applications: We develop a comprehensive section on applications in Operational Research and Management Science, including real-world case studies and interdisciplinary connections with physics.
  6. (6)
    Numerical simulations for $N=1$ and $N=2$: We implement and validate the theory numerically in both one and two spatial dimensions, with detailed interpretation of the results.
1.5. Methodology: The Calculus of Singular Balancing
Our methodology proceeds in four phases, designed to bridge the gap between abstract analysis and practical economic applicability:
Phase 1: Singularity Balance Identification.:
We analyze the HJB equation as a competition between Diffusion ($\Delta u$) and Correction ($b\,h(|\nabla u|)$). By substituting the ansatz $u\sim d^{-\gamma}$, we find the value of $\gamma$ that matches the singular orders. This phase identifies the three asymptotic regimes.
Phase 2: Barrier Construction.:
We use the concave defining function $v(x)$ to construct global sub- and supersolutions that determine the precise boundary blow-up rate. This phase ensures that local boundary estimates propagate correctly to the interior of the domain.
Phase 3: Stochastic Embedding.:
The PDE solution is mapped to a stochastic control problem. The blow-up $u\to\infty$ is interpreted as an impenetrable “cost wall” that enforces the safe-zone boundaries. The Legendre transform mediates the connection between the Hamiltonian and the running cost.
Phase 4: Economic Variable Mapping.:
Mathematical constants are mapped to economic parameters: $a(x)$ to state-dependent discount rates, $b(x)$ to adjustment stiffness coefficients, and $h$ to convex adjustment cost functions. This phase enables direct application to operational research problems.
1.6. Structure of the Paper
The remainder of the paper is organized as follows. Section 2 presents the mathematical framework and all structural assumptions. Section 3 develops the geometric preliminaries and barrier construction. Section 4 provides the complete proof of existence, uniqueness, and convexity (Theorem 2.9 and Theorem 2.11). Section 5 details the proof of sharp boundary asymptotics (Theorem 2.10). Section 6 develops the stochastic control model, the verification theorem, and the full derivation of the Hamilton–Jacobi–Bellman equation. Section 7 presents applications in Operational Research and Management Science, including case studies and interdisciplinary connections with physics. Section 8 describes the monotone iterative scheme. Section 9 presents numerical simulations for both the $N=1$ and $N=2$ cases with detailed interpretation. Section 10 concludes the work. Appendix A and Appendix B contain the complete Python implementation codes for the one-dimensional and two-dimensional solvers, respectively.
2. Mathematical Framework and Assumptions
Throughout this paper, we adopt the following structural and geometric assumptions. These assumptions are designed to be as general as possible while ensuring that the main results (existence, uniqueness, asymptotics, convexity, and stochastic representation) hold simultaneously.
@@ 2. Mathematical Framework and Assumptions
Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) be a bounded open set. The boundary $\partial\Omega$ is of class $C^{2,\alpha}$ for some $\alpha\in(0,1)$. We assume $\Omega$ is strictly convex: there exists a positive constant $\kappa_{0}>0$ such that for every $x\in\partial\Omega$, the principal curvatures $\kappa_{i}(x)$ satisfy:
$\min_{i=1,\dots,N-1}\kappa_{i}(x)\geq\kappa_{0}.$ (2.1)
Remark 2.2 (Geometric significance of strict convexity).
Strict convexity is a natural assumption in both mathematics and economics. In economic dynamics, the domain $\Omega$ represents the set of feasible states (e.g., inventory levels, portfolio weights, production capacities). Strict convexity ensures that any convex combination of two feasible states is also feasible, which is the mathematical formulation of the principle that “diversification reduces risk.” The curvature bound (2.1) ensures that the boundary has no flat segments or corners, which would create singular shocks in the optimal control strategy. In physical terms, the curvature controls the rate at which the distance function $d(x)=\mathrm{dist}(x,\partial\Omega)$ becomes non-smooth, and the lower bound $\kappa_{0}$ guarantees that the tubular neighborhood of $\partial\Omega$ where $d(x)$ is $C^{2}$ has width at least $1/\kappa_{0}$.
@@
$f:\overline{\Omega}\to[0,\infty)$ is a concave function belonging to $C^{k,\alpha}(\overline{\Omega})$ for some $k\geq 1$ and $\alpha\in(0,1)$.
@@
    $h$ is strictly increasing for all $s>0$ and strictly convex;
@@
    There exist constants $l_{1},l_{2}>0$ and $q\in(1,2]$ such that
    $l_{1}\,s^{q}\leq h(s)\leq l_{2}\,s^{q}\quad\text{for all }s\geq 0.$ (2.2)
Remark 2.5 (Economic interpretation of the Hamiltonian).
The Hamiltonian $h$ represents the cost of corrective action per unit of adjustment intensity. The strict convexity (H2) captures the economic principle of increasing marginal costs: the cost of doubling the adjustment rate is more than double the original cost. The power-type growth (H3) with exponent $q\in(1,2]$ encompasses the quadratic case $h(s)=s^{2}$ (commonly used in linear-quadratic control) as well as sub-quadratic costs that arise in models with capacity constraints. The canonical example is $h(s)=s^{q}/q$ for $q\in(1,2]$.
@@
The weights $a,b\in C^{\infty}(\Omega)\cap C^{k,\alpha}(\Omega)$ for $k\geq 1$, are strictly positive in $\Omega$ and satisfy:
@@
    $a_{1}\,d(x)^{\alpha}\leq a(x)\leq a_{2}\,d(x)^{\alpha}$ (2.3)
    near $\partial\Omega$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. We further assume that $a(x)$ is positive convex in $\Omega$.
@@
    There exists a constant $\beta\in\mathbb{R}_{+}=[0,+\infty)$ and limits $b_{1},b_{2}>0$ such that
    $\liminf_{d(x)\to 0}\frac{b(x)}{d(x)^{\beta}}=b_{1}\leq\limsup_{d(x)\to 0}\frac{b(x)}{d(x)^{\beta}}=b_{2}.$ (2.4)
    We assume that $b(x)$ is positive convex in $\Omega$.
Remark 2.7 (Principle of Independence from Data Meaning).
Assumptions 2.12.6 do not assume a specific physical meaning for $x$. This abstraction is intentional: we seek a theory where the boundary dynamics are determined by the category of the singularity (characterized by the exponents $\alpha$, $\beta$, $q$), not by the specific data values. The same equation may describe inventory levels in a warehouse, concentrations in a chemical reactor, or asset allocations in a portfolio, and our results apply identically to all these scenarios.
This structure on the weights mirrors the rigorous bounds employed by Zhang [27] in the context of Monge-Ampère equations with weights, enabling us to derive precise trace limits for our solutions.
Remark 2.8 (Applicability of the Lasry–Lions framework).
The assumptions (H1)–(H3) for the Hamiltonian $h$ together with the strict convexity of $\Omega$ (Assumption 2.1) ensure that the fundamental requirements of the Lasry–Lions framework [17] are fully satisfied. In particular, the strict convexity of $h$ and its power-law growth $h(s)\sim s^{q}$ for $q>1$, combined with the positivity and singular growth of the reaction coefficient $a(x)$, guarantee that the boundary blow-up solution $u$ in (1.3) arises as the value function of a well-posed state-constrained stochastic control problem. Thus, the analytical problem considered in this work is consistently embedded within the classical stochastic paradigm introduced in [17].
2.1. Statement of Main Results
We now state our main contributions precisely.
@@ 2.1. Statement of Main Results
Under Assumptions 2.12.6, the HJB equation (1.3) admits a unique classical solution $u\in C^{2}(\Omega)$ such that $u(x)\to\infty$ as $d(x)\to 0$. In the Gradient-Dominant regime ($q<\beta+2$), the solution satisfies:
$C_{-}\,d(x)^{-\gamma}\leq u(x)\leq C_{+}\,d(x)^{-\gamma}\quad\text{near }\partial\Omega,\quad\gamma=\frac{\beta-q+2}{q-1},$ (2.5)
where $C_{\pm}$ are explicit barrier constants defined in Lemma 3.2.
@@
$\xi_{1}\leq\liminf_{d(x)\to 0}u(x)\,(d(x))^{\gamma}\leq\limsup_{d(x)\to 0}u(x)\,(d(x))^{\gamma}\leq\xi_{2},$ (2.6)
@@
If the limits
$b_{0}=\lim_{d(x)\to 0}b(x)\,d(x)^{-\beta}\quad\text{and}\quad l_{0}=\lim_{s\to\infty}h(s)\,s^{-q}$
exist, then:
$\lim_{d(x)\to 0}u(x)\,(d(x))^{\gamma}=\xi_{0}=\left(\frac{\gamma(\gamma+1)}{b_{0}\,l_{0}\,\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (2.8)
The constant $\xi_{0}$ is the Universal Correction Constant. It represents the precise intensity of the boundary penalty required to enforce the state constraint. Its explicit dependence on $(q,\beta,b_{0},l_{0})$ provides a quantitative tool for calibrating control strategies.
Theorem 2.11 (Convexity Inheritance).
Under Assumptions 2.12.6, if the source term $f(x)$ is concave and the reaction weights $a,b$ are convex in $\Omega$, then the solution $u$ from Theorem 2.9 is strictly convex throughout $\Omega$: $D^{2}u(x)>0$ for all $x\in\Omega$.
Strict convexity is indispensable for practical control: it ensures that the optimal decision for any state $x$ is unique and can be found via simple gradient-descent algorithms. It also guarantees that the running cost functional is jointly convex, which is a prerequisite for the application of duality theory in optimization.